If a company had total sales of $1,000, total variable costs of $400, total fixed costs of $300, and total interest expenses of $100, then the Degree of Financial Leverage would be __.__. Round your answer to 2 decimal places.
Answer #1

Degree of Financial Leverage = EBIT/(EBIT - Interest)

EBIT = sales - variable costs - fixed costs

= 1000-400-300

= 300

DFL = 300/(300-100)

= 300/200

= 1.5